# Efficiently learning depth-3 circuits via quantum agnostic boosting

### v2

@@ A.1 Bond dimension bounds for phase states
where the maximum is over all possible contiguous cuts, call it $L,R$ for left and right and $\textsf{Schmidt-rank}_{L|R}(|\psi\rangle)$ is defined as the Schmidt rank of the state when expressed as $|\psi\rangle=\sumop\displaylimits_{i}c_{i}|\phi_{i}\rangle_{L}\otimes|\gamma_{i}\rangle_{R}$ with $\{|\phi_{i}\rangle_{L}\}_{i},\{|\gamma_{i}\rangle_{R}\}_{i}$ being an orthogonal set of states. Second, since we are dealing with phase states $\frac{1}{\sqrt{2^{n}}}\sumop\displaylimits_{x}(-1)^{f(x)}|x\rangle$ in this work, it is not too hard to see the following: defining $M^{f}_{S}(a,b)=(-1)^{f(a,b)}$ where $a\in{\{0,1\}}^{S},b\in{\{0,1\}}^{\overline{S}}$, then

### v3

@@ A.1 Bond dimension bounds for phase states
where the maximum is over all possible contiguous cuts, call it $L,R$ for left and right and $\textsf{Schmidt-rank}_{L|R}(|\psi\rangle)$ is defined as the Schmidt rank of the state when expressed as $|\psi\rangle=\sumop\displaylimits_{i}c_{i}|\chi_{i}\rangle_{L}\otimes|\gamma_{i}\rangle_{R}$ with $\{|\chi_{i}\rangle_{L}\}_{i},\{|\gamma_{i}\rangle_{R}\}_{i}$ being an orthogonal set of states. Second, since we are dealing with phase states $\frac{1}{\sqrt{2^{n}}}\sumop\displaylimits_{x}(-1)^{f(x)}|x\rangle$ in this work, it is not too hard to see the following: defining $M^{f}_{S}(a,b)=(-1)^{f(a,b)}$ where $a\in{\{0,1\}}^{S},b\in{\{0,1\}}^{\overline{S}}$, then